# DICE: End-to-end Deformation Capture of Hand-Face Interactions from a Single Image

### v1

@@ 4.5 Limitations and Future Works
While our method achieves SotA accuracy on the Decaf [97] dataset and generalizes well to unseen scenes and in-the-wild cases, we still have failure cases when the hand-pose interactions are extremely challenging and have severe occlusions (see Appendix D.2). Moreover, while our method effectively recovers hand and face meshes with visually plausible deformations, there remains room for improvement in deformation accuracy and physical plausibility. In the future, physics-based simulation [37, 56, 38, 31, 60, 41] can be used as a stronger prior, producing more physically accurate estimations. In this paper, although we found using 500 in-the-wild images significantly improves the model’s generalization ability, scaling up to a larger amount of in-the-wild data, on the order of millions or billions, would further enhance performance, which we will study in future work.
@@ 5 Conclusion
In this work, we present DICE, the first end-to-end approach for reconstructing 3D hand and face interaction with deformation from monocular images. Our approach features a two-branch transformer structure, MeshNet and InteractionNet, to model local deform field and global mesh geometry. An inverse-kinematic model, IKNet, is used to output the animatable parametric hand and face meshes. We also proposed a novel weak-supervision training pipeline, using a small amount of in-the-wild images and supervising with a depth prior and an adversarial loss to provide pose priors. Benefitting from our network design and training scheme, DICE demonstrates state-of-the-art accuracy and plausibility, compared with all previous methods. Meanwhile, our method achieves a fast inference speed (20 fps), allowing for more downstream interactive applications. In addition to strong performance on the standard benchmark, DICE also achieves superior generalization performance on in-the-wild data.
@@ A.1 CNN Backbone
The CNN backbone used in our framework is an HRNet-W64 [99], initialized with ImageNet-pretrained weights. The weights of the backbone would be updated during training. We extract a ($49\times H$)-dim feature map from this network and upsamples it to a ($N\times H$)-dim feature map, where $N=N_{h_{k}}+N_{f_{k}}+N_{h_{v}}+N_{h_{v}}$, the total number of head and hand keypoints $N_{h_{k}},N_{f_{k}}$ and vertices $N_{h_{v}},N_{f_{v}}$. Then, we concatenate the keypoints and the vertices corresponding to the head and hand mean pose as keypoints and vertex queries, resulting in a ($(N+3)\times H$)-dim feature map. Random masking of keypoints and vertex queries of rate $30\%$ is applied, following [63].
@@ A.3 IKNet
Our IKNets take in rough mesh features $\mathbf{V}_{F}^{\prime},\mathbf{V}_{H}^{\prime}$ and output the pose and shape parameters $(\theta,\beta)$, as well as the global rotation and translation $(R,T)$. They feature a Multi-Layer Perceptron (MLP) structure, each consisting of five MLP Blocks and a final fully connected layer. Each MLP Block contains a fully connected layer, followed by a batch normalization layer [42] and a ReLU activation layer. There are two skip-connections, connecting the output of the first block with the input of the third block, and the output of the third block with the input of the final fully connected layer. See Fig. 6 for an illustration. The hand and head IKNets have the same structure, differing only in their input and output dimensions. The hidden dimensions of the two IKNets are 1024.
@@ A.4 Training and Testing Details
To be consistent with the training setting of Decaf [^1] Confirmed by the authors of Decaf [97], in the Decaf dataset, we use all eight camera views and the subjects S2, S4, S5, S7, and S8 in the training data split for training. For testing, we use only the front view (view 108) and the subjects S1, S3, and S6 in the testing data split. The low, mid, and high-resolution head mesh consists of $559$, $1675$, and $5023$ vertices, respectively. The low and high-resolution hand mesh consists of $195$ and $778$ vertices, respectively. We use the middle-resolution head mesh and the high-resolution hand mesh as the inputs of head and hand IKNets.
@@ C.1 Mesh losses
$\mathcal{L}_{\text{mesh}}=\mathcal{L}_{\text{reproj}}+4\mathcal{L}_{\text{vert}}+2\mathcal{L}_{\text{key}}+2\mathcal{L}_{\text{params}}.$ (6)
Vertices Loss $L_{1}$ loss is used for predicted rough 3D face and hand vertices ${\mathbf{V}_{f}^{\prime}}$, ${\mathbf{V}_{h}^{\prime}}$, FLAME-regressed undeformed 3D face vertices ${{\mathbf{V}_{f}^{*}}}$ and MANO-regressed 3D hand vertices ${{\mathbf{V}_{h}}}$ against the ground-truth 3D undeformed face vertices $\mathbf{\hat{V}}_{f}$ and 3D hand vertices $\mathbf{\hat{V}}_{h}$.
$\mathcal{L}_{\text{vert}}=\lambda_{h}(\mu_{\text{nonpara}}\|\mathbf{V}_{h}^{\prime}-\mathbf{\hat{V}}_{h}\|_{1}+\|\mathbf{V}_{h}-\mathbf{\hat{V}}_{h}\|_{1})+\lambda_{f}(\mu_{\text{nonpara}}\|\mathbf{V}_{f}^{\prime}-\mathbf{\hat{V}}_{h}\|_{1}+\|\mathbf{V}_{f}^{*}-\mathbf{\hat{V}}_{f}\|_{1}),$ (7)
@@
Keypoints Loss We use $L_{1}$ loss for predicted rough 3D face and hand keypoints $\mathbf{K}_{f}^{\prime}$, $\mathbf{K}_{h}^{\prime}$, 3D face and hand keypoints extracted from rough mesh ${\mathbf{K}_{f_{\text{mesh}}}},{\mathbf{K}_{h_{\text{mesh}}}}$, FLAME-regressed 3D face keypoints ${\mathbf{K}_{f}}$ and MANO-regressed 3D hand keypoints ${\mathbf{K}_{h}}$ against the ground-truth 3D undeformed face keypoints ${\mathbf{\hat{K}}_{f}}$ and 3D hand keypoints ${\mathbf{\hat{K}}_{f}}$.
@@
$+\|{\mathbf{K}_{f}}-{\mathbf{\hat{K}}_{f}}\|_{1}+\|{\mathbf{K}_{h}}-{\mathbf{\hat{K}}_{h}}\|_{1},$ (9)
@@
Reprojection loss $L_{1}$ loss is used for reprojected rough 3D face and hand keypoints $\mathbf{K}_{f}^{\prime}$, $\mathbf{K}_{h}^{\prime}$, 3D face and hand keypoints extracted from rough mesh ${\mathbf{K}_{f_{\text{mesh}}}},{\mathbf{K}_{h_{\text{mesh}}}}$, FLAME-regressed 3D face keypoints ${\mathbf{\hat{K}}_{f}}$ and MANO-regressed 3D hand keypoints ${\mathbf{\hat{K}}_{h}}$ against the ground-truth face and hand 2D keypoints $\mathbf{\hat{K}}_{f_{\text{2D}}},\mathbf{\hat{K}}_{h_{\text{2D}}}$.
@@
$+\lambda_{f}(\|\Pi(\mathbf{K}_{f}^{\prime})-\mathbf{\hat{K}}_{f_{\text{2D}}}\|_{1}+\|\Pi({\mathbf{K}_{f_{\text{mesh}}}})-\mathbf{\hat{K}}_{f_{\text{2D}}}\|_{1}+\|\Pi({\mathbf{K}_{f}})-\mathbf{\hat{K}}_{f_{\text{2D}}}\|_{1}),$ (11)
@@
Parameter loss We apply $L_{1}$ loss on the regressed hand and face pose, shape, and facial expression parameters against their respective ground truths.
@@
C.2 Interaction losses
The interaction loss $\mathcal{L}_{\text{interaction}}$ consists of four components:
$\mathcal{L}_{\text{interaction}}=0.2\mathcal{L}_{\text{touch}}+0.6\mathcal{L}_{\text{contact}}+\mathcal{L}_{\text{collision}}+6\mathcal{L}_{\text{deform}}.$ (15)
Deformation loss Due to the human anatomy, some vertices on the face are more easily deformed than other vertices. Therefore, we impose an adaptive weighting on each vertex and use square loss to penalize large deformation. We also have a regularization term to penalize extremely large deformations.
$\mathcal{L}_{\text{deform}}=\sum_{i\in\mathcal{I}}(1+\mu\|\hat{d_{i}}\|_{2})\|\hat{d_{i}}-d_{i}\|_{2}^{2}+\lambda\sum_{i\in\mathcal{L}}\|d_{i}\|,$ (16)
@@ C.2 Interaction losses
Touch loss Let $\mathbf{V}_{F_{C}}$ and $\mathbf{V}_{H_{C}}$ denote the set of face and hand vertices that are predicted by the model to have contact probability greater than $0.5$.
$\mathcal{L}_{\text{touch}}=\text{CD}(\mathbf{V}_{F_{C}},\mathbf{V}_{H_{C}})+\text{CD}(\mathbf{V}_{H_{C}},\mathbf{V}_{F_{C}}),$ (17)
@@
Collision loss Let $\mathbf{V}_{H_{\text{Col}}}$ denote the set of hand vertices that penetrates the face surface, $\mathbf{V}_{F}$ and $\mathbf{D}_{F}$ denote the predicted face mesh vertices and deformations.
$\mathcal{L}_{\text{collision}}=\text{CD}(\mathbf{V}_{H_{\text{Col}}},\mathbf{V}_{F}-\mathbf{D}_{F}).$ (18)
Contact loss Let $\mathbf{C}_{H}$ and $\mathbf{C}_{F}$ denote the predicted hand and face contact probabilities and ${\mathbf{\hat{C}}_{H}}$, ${\mathbf{\hat{C}}_{F}}$ denote the ground-truth contact labels.
$\mathcal{L}_{\text{contact}}=\text{BCE}(\mathbf{C}_{H},{\mathbf{\hat{C}}_{H}})+\text{BCE}({\mathbf{C}_{F}},{\mathbf{\hat{C}}_{F}}),$ (19)
where BCE denotes the binary cross-entropy loss.
@@ Appendix D More Discussions
Figure 8: Examples of failed keypoint estimation in case of large self-occlusion. (a) input image; (b) inaccurate keypoint estimation by the same keypoint estimators used in Decaf [68, 3]; (c) reconstructed hand-face interaction by our method. (d) reconstructed hand-face interaction by Decaf.
As seen in Fig. 8, our end-to-end DICE method is robust under challenging self-occlusion cases, such as the hand covering more than half of the face. On the other hand, Decaf [97], which requires an initial keypoint prediction for test-time optimization, performs poorly in this situation.
@@ D.2 Failure Cases
In Fig. 9, we demonstrate the failure cases of our method. As shown in Fig. 9 (a), when there is a complex interaction between the hand and face, such as the presence of a cleaning sponge, there is a drop in the reconstruction accuracy of the hand mesh recovery. Also, as in Fig. 9 (b), When the face completely occludes the hand, a highly challenging scenario unseen in the training data, our model could not faithfully reconstruct the hand position.
@@
Figure 9: Samples of failure cases.

### v2

@@ 4.5 Limitations and Future Works
While our method achieves SotA accuracy on the Decaf (Shimada et al., 2023) dataset and generalizes well to unseen scenes and in-the-wild cases, it still encounters failure cases when the hand-pose interactions are extremely challenging and have severe occlusions (see Appendix D.2). Moreover, despite our method effectively recovering hand and face meshes with visually plausible face deformations, there remains room for improvement in deformation accuracy and physical plausibility. Hand deformations could also be considered in future work for more realistic reconstructions. In the future, physics-based simulation (Hu et al., 2018; Li et al., 2020; Hu et al., 2019; Han et al., 2019; Lin et al., 2022; Huang et al., 2024) can be used as a stronger prior, producing more physically accurate estimations. In this paper, although we found using 500 in-the-wild images significantly improves the model’s generalization ability, scaling up to a larger amount of in-the-wild data, on the order of millions or billions, would further enhance performance, which we will study in future work.
@@ 5 Conclusion
In this work, we present DICE, the first end-to-end approach for reconstructing 3D hand and face interaction with deformation from monocular images. Our approach features a two-branch transformer structure, MeshNet, and InteractionNet, to model local deform field and global mesh geometry. An inverse-kinematic model, IKNet, is used to output the animatable parametric hand and face meshes. We also proposed a novel weak-supervision training pipeline, using a small amount of in-the-wild images and supervising with a depth prior and an adversarial loss to provide pose priors. Benefitting from our network design and training scheme, DICE demonstrates state-of-the-art accuracy and plausibility, compared with all previous methods. Meanwhile, our method achieves a fast inference speed (20 fps), allowing for more downstream interactive applications. In addition to strong performance on the standard benchmark, DICE also achieves superior generalization performance on in-the-wild data.
Acknowledgments
This work is partly supported by the Research Grant Council of Hong Kong (Ref: 17210222), the Innovation and Technology Commission of Hong Kong under the ITSP-Platform grants (Ref: ITS/319/21FP, ITS/335/23FP) and the InnoHK initiative (TransGP project). The research work was in part conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust.
@@ A.1 CNN Backbone
The CNN backbone used in our framework is an HRNet-W64 (Sun et al., 2019), initialized with ImageNet-pretrained weights. The weights of the backbone would be updated during training. We extract a ($49\times H$)-dim feature map from this network and upsamples it to a ($N\times H$)-dim feature map, where $N=N_{h_{k}}+N_{f_{k}}+N_{h_{v}}+N_{h_{v}}$, the total number of head and hand keypoints $N_{h_{k}},N_{f_{k}}$ and vertices $N_{h_{v}},N_{f_{v}}$. Then, we concatenate the keypoints and the vertices corresponding to the head and hand mean pose as keypoints and vertex queries, resulting in a ($(N+3)\times H$)-dim feature map. Random masking of keypoints and vertex queries of rate $30\%$ is applied, following (Lin et al., 2021a).
@@ A.3 IKNet
Our IKNets take in rough mesh features $\mathbf{V}_{F}^{\prime},\mathbf{V}_{H}^{\prime}$ and output the pose and shape parameters $(\theta,\beta)$, as well as the global rotation and translation $(R,T)$. They feature a Multi-Layer Perceptron (MLP) structure, each consisting of five MLP Blocks and a final fully connected layer. Each MLP Block contains a fully connected layer, followed by a batch normalization layer (Ioffe & Szegedy, 2015) and a ReLU activation layer. There are two skip-connections, connecting the output of the first block with the input of the third block, and the output of the third block with the input of the final fully connected layer. See Fig. 6 for an illustration. The hand and head IKNets have the same structure, differing only in their input and output dimensions. The hidden dimensions of the two IKNets are 1024.
@@ A.4 Training and Testing Details
To be consistent with the training setting of Decaf [^1] Confirmed by the authors of Decaf (Shimada et al., 2023), in the Decaf dataset, we use all eight camera views and the subjects S2, S4, S5, S7, and S8 in the training data split for training. For testing, we use only the front view (view 108) and the subjects S1, S3, and S6 in the testing data split. The low, mid, and high-resolution head mesh consists of $559$, $1675$, and $5023$ vertices, respectively. The low and high-resolution hand mesh consists of $195$ and $778$ vertices, respectively. We use the middle-resolution head mesh and the high-resolution hand mesh as the inputs of head and hand IKNets.
@@ C.1 Mesh losses
$\mathcal{L}_{\text{mesh}}=\mathcal{L}_{\text{reproj}}+4\mathcal{L}_{\text{vert}}+2\mathcal{L}_{\text{key}}+2\mathcal{L}_{\text{params}}$ (6)
Vertices Loss. $L_{1}$ loss is used for predicted rough 3D face and hand vertices ${\mathbf{V}_{f}^{\prime}}$, ${\mathbf{V}_{h}^{\prime}}$, FLAME-regressed undeformed 3D face vertices ${{\mathbf{V}_{f}^{*}}}$ and MANO-regressed 3D hand vertices ${{\mathbf{V}_{h}}}$ against the ground-truth 3D undeformed face vertices $\mathbf{\hat{V}}_{f}$ and 3D hand vertices $\mathbf{\hat{V}}_{h}$.
$\mathcal{L}_{\text{vert}}=\lambda_{h}(\mu_{\text{nonpara}}\|\mathbf{V}_{h}^{\prime}-\mathbf{\hat{V}}_{h}\|_{1}+\|\mathbf{V}_{h}-\mathbf{\hat{V}}_{h}\|_{1})+\lambda_{f}(\mu_{\text{nonpara}}\|\mathbf{V}_{f}^{\prime}-\mathbf{\hat{V}}_{h}\|_{1}+\|\mathbf{V}_{f}^{*}-\mathbf{\hat{V}}_{f}\|_{1})$ (7)
@@
Keypoints Loss. We use $L_{1}$ loss for predicted rough 3D face and hand keypoints $\mathbf{K}_{f}^{\prime}$, $\mathbf{K}_{h}^{\prime}$, 3D face and hand keypoints extracted from rough mesh ${\mathbf{K}_{f_{\text{mesh}}}},{\mathbf{K}_{h_{\text{mesh}}}}$, FLAME-regressed 3D face keypoints ${\mathbf{K}_{f}}$ and MANO-regressed 3D hand keypoints ${\mathbf{K}_{h}}$ against the ground-truth 3D undeformed face keypoints ${\mathbf{\hat{K}}_{f}}$ and 3D hand keypoints ${\mathbf{\hat{K}}_{f}}$.
@@
$+\|{\mathbf{K}_{f}}-{\mathbf{\hat{K}}_{f}}\|_{1}+\|{\mathbf{K}_{h}}-{\mathbf{\hat{K}}_{h}}\|_{1}$ (9)
@@
Reprojection loss. $L_{1}$ loss is used for reprojected rough 3D face and hand keypoints $\mathbf{K}_{f}^{\prime}$, $\mathbf{K}_{h}^{\prime}$, 3D face and hand keypoints extracted from rough mesh ${\mathbf{K}_{f_{\text{mesh}}}},{\mathbf{K}_{h_{\text{mesh}}}}$, FLAME-regressed 3D face keypoints ${\mathbf{\hat{K}}_{f}}$ and MANO-regressed 3D hand keypoints ${\mathbf{\hat{K}}_{h}}$ against the ground-truth face and hand 2D keypoints $\mathbf{\hat{K}}_{f_{\text{2D}}},\mathbf{\hat{K}}_{h_{\text{2D}}}$.
@@
$+\lambda_{f}(\|\Pi(\mathbf{K}_{f}^{\prime})-\mathbf{\hat{K}}_{f_{\text{2D}}}\|_{1}+\|\Pi({\mathbf{K}_{f_{\text{mesh}}}})-\mathbf{\hat{K}}_{f_{\text{2D}}}\|_{1}+\|\Pi({\mathbf{K}_{f}})-\mathbf{\hat{K}}_{f_{\text{2D}}}\|_{1})$ (11)
@@
Parameter loss. We apply $L_{1}$ loss on the regressed hand and face pose, shape, and facial expression parameters against their respective ground truths.
@@
C.2 Interaction losses.
The interaction loss $\mathcal{L}_{\text{interaction}}$ consists of four components.
$\mathcal{L}_{\text{interaction}}=0.2\mathcal{L}_{\text{touch}}+0.6\mathcal{L}_{\text{contact}}+\mathcal{L}_{\text{collision}}+6\mathcal{L}_{\text{deform}}$ (15)
Deformation loss. Due to the human anatomy, some vertices on the face are more easily deformed than other vertices. Therefore, we impose an adaptive weighting on each vertex, and use square loss to penalize large deformation. We also have a regularization term to penalize extremely large deformations.
$\mathcal{L}_{\text{deform}}=\sum_{i\in\mathcal{I}}(1+\mu\|\hat{d_{i}}\|_{2})\|\hat{d_{i}}-d_{i}\|_{2}^{2}+\lambda\sum_{i\in\mathcal{L}}\|d_{i}\|$ (16)
@@
Touch loss. Let $\mathbf{V}_{F_{C}}$ and $\mathbf{V}_{H_{C}}$ denote the set of face and hand vertices that are predicted by the model to have contact probability greater than $0.5$.
$\mathcal{L}_{\text{touch}}=\text{CD}(\mathbf{V}_{F_{C}},\mathbf{V}_{H_{C}})+\text{CD}(\mathbf{V}_{H_{C}},\mathbf{V}_{F_{C}})$ (17)
@@
Collision loss. Let $\mathbf{V}_{H_{\text{Col}}}$ denote the set of hand vertices that penetrates the face surface, $\mathbf{V}_{F}$ and $\mathbf{D}_{F}$ denote the predicted face mesh vertices and deformations.
$\mathcal{L}_{\text{collision}}=\text{CD}(\mathbf{V}_{H_{\text{Col}}},\mathbf{V}_{F}-\mathbf{D}_{F})$ (18)
Contact loss. Let $\mathbf{C}_{H}$ and $\mathbf{C}_{F}$ denote the predicted hand and face contact probabilities and ${\mathbf{\hat{C}}_{H}}$, ${\mathbf{\hat{C}}_{F}}$ denote the ground-truth contact labels.
$\mathcal{L}_{\text{contact}}=\text{BCE}(\mathbf{C}_{H},{\mathbf{\hat{C}}_{H}})+\text{BCE}({\mathbf{C}_{F}},{\mathbf{\hat{C}}_{F}})$ (19)
Where BCE denote the binary cross-entropy loss.
@@ Appendix D More Discussions
Figure 8: Examples of failed keypoint estimation in case of large self-occlusion. (a) input image; (b) inaccurate keypoint estimation by the same keypoint estimators used in Decaf (Lugaresi et al., 2019; Bulat & Tzimiropoulos, 2017); (c) reconstructed hand-face interaction by our method. (d) reconstructed hand-face interaction by Decaf.
As seen in Fig. 8, our end-to-end DICE method is robust under challenging self-occlusion cases, such as the hand covering more than half of the face. On the other hand, Decaf (Shimada et al., 2023), which requires an initial keypoint prediction for test-time optimization, performs poorly in this situation.
@@ D.2 Failure Cases
In Fig. 9, we demonstrate the failure cases of our method. When the hand is extremely far from the face, or when the hand is completely obscured by the head, our method could fail to reconstruct the hand-face interaction. Also, when given out-of-distribution data, such as when the hand is wearing gloves or the input subject is an infant, the reconstruction accuracy could degrade.
@@
Figure 9: Examples of failure cases in case of complete occlusion of the hand.
(a) Hand or face completely occluded. (b) Out-of-distribution data.
D.3 Societal Impact
D.3.1 Potential Misuse
DICE enables tracking of individuals’ appearances, gestures, and interactions with high fidelity, there is a risk that it may be misused for negative applications, such as surveillance, and may cause privacy infringement. Also, since DICE makes use of a readily animatable representation, it could enable realistic deepfakes driven by the pose and shape information collected, which could be used in creating misinformation and conducting identity theft. We are firmly against any form of misuse of the DICE model.
D.3.2 Data Fairness
As hand-face interaction recovery is a human-related task, data fairness is critical. The currently used Decaf Shimada et al. (2023) dataset needs improvement in the inclusion of human actors from underrepresented demographic groups. This may result in a model trained only on Decaf underperforming on input data on such groups, perpetuating inequality and limiting equitable access. Our weak-supervised training scheme introduces diverse in-the-wild data, which could alleviate this issue as the amount of in-the-wild data scales up.